Matthew is 3 times as old as Jenny. In 7 years he will be twice as old as she will be then. How old is each now.
I have figured out the answer, but not how to work the problem. I tried:
x + 1/3X = 2 times X + 7 + 1/3 X + 7 which is completely wrong. I have not done a problen where I didn't have a specific number for an answer. Reid

Found 3 solutions by scott8148, ankor@dixie-net.com, oberobic:
Answer by scott8148(6628)

M = 3J

M + 7 = 2(J + 7)

Answer by ankor@dixie-net.com(22740)

Matthew is 3 times as old as Jenny.
m = 3j
In 7 years he will be twice as old as she will be then.
m + 7 = 2(j+7)
m + 7 = 2j + 14
m = 2j + 14 - 7
m = 2j + 7
Replace m with 3j
3j = 2j + 7
3j - 2j = 7
j = 7 yrs
then obviously
m = 21 yrs



Answer by oberobic(2304)

M = 3*J
.
In 7 years, their ages will be:
M+7
J+7
.
M+7 = 2*(J+7)
.
Substitute M=3*J
.
3J + 7 = 2J + 14
.
J = 7
.
M = 3J = 21
.
Check their age relationship in 7 years.
M+7 = 28
J+7= 14
M+7 = 2(J+7)
Correct.
.
Answer: Matthew is 21, and Jenny is 7.
.
Done.